Solve the following equation for the unknown:
1) 6.2(x - 15.1) = 9.3(-x - 3.2)

Answer by checkley77(12569)

6.2(x - 15.1) = 9.3(-x - 3.2)
6.2x-93.62=-9.3x-29.76
6.2x+9.3x=93.62-29.76
15.5x=63.86
x=63.86/15.5
x=4.12 answer.
proof:
6.2(4.12-15.1)=9.3(-4.12-3.2)
6.2*-10.98=9.3*7.32
68.076=68.076